OpenStudy (helpmeee123):

Estimate to the nearest integer Square root of 7 .

OpenStudy (unklerhaukus):

\[\sqrt4<\sqrt 7<\sqrt 9\]

OpenStudy (helpmeee123):

What is tbeanswer?

OpenStudy (anonymous):

\[\sqrt{7}=2.64575131106\] so round this to the nearest integer

OpenStudy (anonymous):

so what i did was I went to the calculator and I put square root of 7 and that answer came up THATS IT NOTHING ELSE NO SECRET

OpenStudy (helpmeee123):

Ahhhh, okay! But without using calculator how will we know?